the following first order reaction has a half life of 22.5 minutes.

2H2O2(aq)\rightarrow 2H2O(l) + O2(g)

a. calculate the rate constant for this reaction.

b. if the initial concentraition of H2O2 is 0.30M, calculate the amout of H2O2 that reamains after 45 min.
Answer #1

a) we know that the relation between rate constant and half life is:

K = t1/2 / 0.693

t1/2 = 22.5 minutes

K = 22.5 / 0.693 = 32.46 min-1

b) The half lie is 22.5 minutes

If initial concentration of H2O2 = 0.30 M

so after 22.5 minutes the concentratio of H2O2 left = 0.15 M

So after one more half life ( total time = 45min), concentration of H2O2 left = 0.15 /2   = 0.075 M